OpenStudy (anonymous):

Which statement is true about both solutions? *Equation 1 |5x-6|=-41 *Equation 2 |7x+13|=27 A. Equation 1 has more solutions than equation 2 B. Equation 1 & 2 have the same # of solutions. C. Equation 2 has more solutions than equation 1 D. The number of solutions cannot be determined

OpenStudy (anonymous):

a

OpenStudy (anonymous):

equation 1 has no solution at all

OpenStudy (anonymous):

you have to remove mode only

OpenStudy (anonymous):

i have no idea how modulus of something can be negative

OpenStudy (anonymous):

this |5x-6|=-41 open one time without change in sign and one time with whole -

OpenStudy (anonymous):

we get one time 5x - 6 = -41 and one time -5x + 6 = 41 then solve both equation